OpenStudy (anonymous):

Find the discriminant and describe the nature of the roots. a. 3x^2+7x-10=0 b.5x^2-4x+8=0 c.36^2+22x+25=0 d.7x^2-2x-6=0

OpenStudy (anonymous):

for ax2+bx+c = 0 discriminant=b^2 -4ac a) D= 7^2 -4(3)(-10) =49+120=169(real positive perfect square nence roots are real unequal and rational)

OpenStudy (anonymous):

ohhh, I get it now.. Thank you!

OpenStudy (anonymous):

welcome...